OpenStudy (anonymous):

When will 30x(1.2^x)=45+3x

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Ohh its not 30x it's just 30

OpenStudy (anonymous):

is there answer choices?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It's hard

OpenStudy (anonymous):

oh okay. and yeah i have been studying these in class but i forgot.:/

OpenStudy (anonymous):

Ohh ok

OpenStudy (anonymous):

Me too

OpenStudy (anonymous):

Oh. Sorry I wasn't much help.

OpenStudy (anonymous):

It's ok

OpenStudy (xapproachesinfinity):

is that \[30(1.2^x)=45+3x\]

OpenStudy (anonymous):

Yeah

OpenStudy (xapproachesinfinity):

So do you think you need to do?

OpenStudy (solomonzelman):

Well, I would graph the two functions (on desmos.com for example), and see where they interect.

OpenStudy (solomonzelman):

(where - at which value of x)

OpenStudy (xapproachesinfinity):

yeah! but there is an algebraic way

OpenStudy (xapproachesinfinity):

i already looked at the graphs two solution exist for this

OpenStudy (anonymous):

xapproachesinfinity is right. In the situation that no calculators are allowed and this is a test question, or in the situation that you want to algebraically check your work, knowing how to do the algebra on this one can go a long way... :D

OpenStudy (anonymous):

If you are using the graph method, and there are TWO solutions, you will need to plug them in each and see which value makes a true statement. ^^

OpenStudy (xapproachesinfinity):

let's how it is done algebraically

OpenStudy (xapproachesinfinity):

i would say algebra alone won't do! then let's see what you know? have you taken calculus

OpenStudy (xapproachesinfinity):

so what do you say? i need to confirm what you know with calculus we can figure out this problem

OpenStudy (xapproachesinfinity):

we just need to let \[f(x)=10(1.2)^x -x-15\]